OpenStudy (anonymous):

Will give Medals and Fan!!!! Use the substitution method to solve the following systems of equations. Write the solution as an ordered pair, and show all work. x+y=5 and y=x-7

OpenStudy (anonymous):

Quick answers please!

OpenStudy (unklerhaukus):

subsiitute the second equation y=x-7 into the first x+ y =5 x+(x-7) =5 solve this for x

OpenStudy (anonymous):

Thanks!

OpenStudy (unklerhaukus):

what do you get?

OpenStudy (unklerhaukus):

....

OpenStudy (unklerhaukus):

after you find x , back-substitute to find y

OpenStudy (unklerhaukus):

@iCreeperrr have you worked out x?

OpenStudy (anonymous):

no... I forgot how xD

OpenStudy (anonymous):

it would be simplified to x-7x=5 right?

OpenStudy (unklerhaukus):

x+y = 5 and y = x-7 x+(x-7) = 5 x+x-7 = 5

OpenStudy (unklerhaukus):

what is x+x?

OpenStudy (anonymous):

x^2

OpenStudy (unklerhaukus):

nope x times x is x squared , by x + x = (1+1)x =

OpenStudy (anonymous):

2x.

OpenStudy (unklerhaukus):

that's better

OpenStudy (unklerhaukus):

so you have x+(x-7) = 5 x+x-7 = 5 2x - 7 =5 now add seven to both sides 2x - 7 + 7 = 5+ 7 what does this simplify to ?

OpenStudy (anonymous):

x+(x-7)=5 x+x-7=5 2x-7=5 2x=12 x=6 right?

OpenStudy (unklerhaukus):

thats right you have found x, now we can find y

OpenStudy (anonymous):

then substitute that into the other equation

OpenStudy (anonymous):

y=6-7 y=1

OpenStudy (unklerhaukus):

remember the second equation was y = x-7 so y = 6 -7

OpenStudy (unklerhaukus):

almost, but 6-7 is not equal to 1

OpenStudy (anonymous):

-1 I meant xD

OpenStudy (unklerhaukus):

haha thats better,

OpenStudy (anonymous):

Thanks for the help, its been a while since I learned this and I don't have my book with me.. :l

OpenStudy (unklerhaukus):

so you have x=6 and y=-1 if we check these in the original equations they should work x+y=5 and y=x-7 6+ -1 = 5 ? -1 = 6-7 ? if theses are both true we have correctly solved for x and y

OpenStudy (anonymous):

That's what I got :)

OpenStudy (unklerhaukus):

goood

OpenStudy (unklerhaukus):

now to write the solution as an ordered pair just write the solution in the form (x, y)

OpenStudy (anonymous):

(6,-1)

OpenStudy (unklerhaukus):